### v1

@@ 7 Conclusion
Prompt-based in-context learning is an efficient machine learning method, requiring no parameter updates to the LLM. While achieving good performance on language tasks (System 1), performance on reasoning tasks (System 2) was lacking. Reasoning tasks, such as math word problems, are typically solved in step-by-step fashion. Recently prompts have been developed that guide an LLM to “think step by step” (Chain-of-thought), and to evaluate and verify the step results. The performance of reasoning with LLMs has improved greatly. Together, the surveyed methods allow the LLM to follow high-quality multi-step reasoning chains. Python code or other formal languages have been used successfully to reduce the error in reasoning steps. Also, in the field of autonomous agents and robotic action, good performance has been achieved by grounding reasoning answers in the environment and the physical constraints of robotic movement.
@@
For complex reasoning tasks a large number of reasoning steps may be generated. To control the size of the reasoning space interactively, external scaffolding algorithms can be used. Often, variations on search algorithms or reinforcement learning are used. The symbolic and connectionist AI traditions come together in reasoning prompts and search algorithms that help LLM neural networks solve natural language math word and related problems.
@@
Among the most popular reasoning benchmarks in this survey is GSM8K, which contains 8500 grade school math word problems. With LLMs such as GPT-3, reasoning approaches show an improvement of 20-50% points over standard prompting methods. For further progress in the field, the development of other challenging benchmarks is important.
@@
The field of reasoning with LLMs is quite new, and theoretical understanding is lacking in important areas, such as faithful reasoning (models may sometimes find the right answer for the wrong reason). Although prompt-based learning allows few-shot learning, the computational needs of LLMs pretraining and finetuning are still high, hence the interest in small language models. Reasoning skills that work in large models can often be transferred to small models.
@@
LLM-reasoning is an active field of research, with connections to artificial general intelligence. The field has shown great progress. Based on current limitations and open questions we provide a research agenda highlighting opportunities for further progress in harder reasoning problems, metacognition, and small language models, amongst others.

### v2

@@ 5. Conclusion
Prompt-based in-context learning is an efficient machine learning method, requiring no parameter updates to the LLM. While achieving good performance on language tasks, performance on reasoning tasks was lacking originally. Reasoning tasks, such as math word problems, are typically solved in a step-by-step fashion. Recently prompts have been developed that guide an LLM to “think step by step” (Chain-of-thought), and to evaluate and verify the step results. The performance of multi-step reasoning with LLMs has improved greatly, and the field has progressed beyond math word problems. Together, the surveyed methods allow the LLM to follow high-quality multi-step reasoning chains. Python code or other formal languages have been used successfully to reduce the error in reasoning steps. Also, in the field of autonomous agents and robotic action, good performance has been achieved by grounding reasoning answers in the environment and the physical constraints of robotic movement.
@@
For complex reasoning tasks a large number of reasoning steps may be generated. To control the size of the multi-step reasoning space dynamically, external scaffolding algorithms can be used. Often, variations on search algorithms are used, and especially reinforcement learning. The symbolic and connectionist AI traditions come together in reasoning prompts and search algorithms that help LLM neural networks solve natural language math word and related problems. Inference-time reasoning results can be used to augment finetuning, in a feedback loop.
LLMs hallucinate and suffer from bias, and their use poses different ethical dangers. In multi-step reasoning methods, self-verification methods have been developed to reduce error-accumulation, yet ethical dangers may remain, especially since reflection is not error-free. Additionally, retrieval augmentation methods ground LLM output directly in sources such as Wikipedia.
@@
Among the most popular reasoning benchmarks in this survey is GSM8K, which contains 8500 grade school math word problems. With older LLMs such as GPT-3, reasoning approaches show an improvement of 20-50% points over standard prompting methods, and some even more. Many current LLMs have since adopted the reasoning approaches. The success of reasoning with LLMs has attracted more applications, and with them, benchmarks are diverging.
@@
The field of reasoning with LLMs is quite new, and theoretical understanding is lacking in important areas, such as faithful reasoning (models may sometimes find the right answer using an incorrect reasoning chain). There may be more opportunities for efficiency gains by integrating the training of reasoning models, using reinforcement learning, both in training and in context. Although prompt-based learning allows few-shot learning at inference time, the computational needs of LLM pretraining and finetuning are still high, hence the interest in small language models. Reasoning skills that work in large models can often be distilled to small models.
@@
LLM-reasoning is an active field of research that shows great progress. Based on current limitations and open questions we provide a research agenda highlighting opportunities for further progress with reinforcement learning (for finetuning, in context reinforcement learning and self-reflection), for LLMs that generate code for external tools, for multimodal world models, for value-aligned reasoning, and for small language models, among others.
Acknowledgements
We thank the anonymous reviewers for their valuable suggestions that have considerably improved the paper.